Consider the following quadratic equation: 5x^2-37x+17=3
*Note: Solve by factoring.

Answer by lwsshak3(11628)

Consider the following quadratic equation:
5x^2-37x+17=3
5x^2-37x+14=0
(5x-2)(x-7)=0
x=2/5
or
x=7